# Gravity-Informed Deep Learning Framework for Predicting Ship Traffic Flow and Invasion Risk of Non-Indigenous Species via Ballast Water Discharge

### v1

@@ Results
After analyzing the results from the model shown in Figure 3(a), we observed that the Random Forest model has a higher accuracy. However, the Logistic Regression model acts as a filter by removing some 1- and 2-trip trajectories before predicting ship traffic flow. These trajectories are considered unstable and unreliable for predicting future mobility flow and may be outliers. Therefore, the Logistic Regression model is used to identify and remove these trajectories, leaving behind only those predicted to have more stable traffic flow between origin-destination pairs. These trajectories are then utilized as features of gravity-informed models for improving ship traffic flow prediction, such as seen in Figure 5, which allows for the environmental similarity analysis and the ballast water risk assessment discussed as follows.
Figure 5: Experimental pipeline for predicting ship traffic flows with gravity-informed models and assessing environmental similarity for ballast water risk assessment (continuation from Figure 3).
Gravity Model
As in the laws of gravity, the gravity model describes the interaction between two entities proportionally to their masses and inversely proportional to their distance [22]. Over time, the model has been adapted to solve various practical problems, such as application in studying human mobility and migration patterns [23, 24], which follows the gravity basis that the number of commuters between two locations is related to the populations of the two locations and the physical distances. In addition, the gravity model also includes adjustable parameters that can be learned from historical data. Other popular applications include forecasting economic interactions [27], communication networks [28], and cargo shipping [15, 50, 29].
Even though the gravity model has prevailed for many years, its limitations can hardly be overlooked. First, the simple components of “masses” and “distances” cannot capture the connection with real mobility flows and lack comprehensive factors that can represent specific scenarios [32], such as the city infrastructure in urban mobility studies and the global trade pattern impacting cargo shipping flows. Also, the lack of limits on the flow increase can lead to the predicted flow size larger than the source “masses”, making the model predictions unreasonable and challenging to interpret [32, 42]. Additionally, it is difficult to deal with multiple adjustable parameters in the model for the prediction without enough previously observed data.
Alternative solutions have been proposed to overcome the limitations of gravity models in predicting mobility flows. In this sense, the radiation model offers a parameter-free approach which resolves the problem of having multiple parameters in the gravity model. Besides, it limits the number of flows by introducing the total number of individuals departing from the source location [32, 34]. Moreover, different studies used multiple relative factors, such as employment and urban infrastructure, that capture more important characteristics to improve prediction accuracy and adaptability of the model [42, 41]. The evolving nature of artificial intelligence and the rise of large language models have brought about new technologies. Therefore, improving the state of the art by using more capable technologies over new arrangements and combinations of data is essential. This becomes more evident when considering evolving factors such as climate change and recurrent anthropogenic effects observed on the oceans.
Traditional Deep Gravity
Deep Gravity is a recent study in urban mobility that integrates the gravity model with deep neural networks [41]. Unlike traditional gravity models that rely on maximum likelihood estimation for parameter tuning, Deep Gravity utilizes a cross-entropy loss function as a substitute for maximum likelihood estimation. This innovative approach effectively adapts the gravity model to a neural network framework. Also, it departs from the traditional multi-parameter structure of gravity models; instead, it incorporates the populations, distances, and infrastructures into the feature set and integrates data from various sources.
However, the simplicity of the multilayer perceptrons (MLPs) structure in Deep Gravity presents certain limitations. Specifically, the complex multivariate relationships inherent in mobility flows can be challenging to capture accurately through the composition of multiple linear functions. Moreover, the deep architecture composed of fifteen linear layers stacked in sequence demands many model parameters. Thus, it requires more computational resources and training time to provide a viable solution. While Deep Gravity utilizes a multilayer architecture, the underlying gravity model can be simplified to contain one or a couple of layers. In this study, we employed multiple multi-layered models as benchmarks for comparison against our proposed framework that leverages Transformers on Deep Gravity Models.
Transformer Deep Gravity
In this study, we incorporate the self-attention mechanism of the Transformer architecture [44] into our proposed framework. Compared to the conventional MLPs structure, the self-attention mechanism can inspect the input sequence and weigh to identify and prioritize the most relevant elements for generating the output, and this characteristic enables our model to capture crucial dependencies in vessel mobility flows effectively. Additionally, the self-attention mechanism accomplishes high performance with fewer parameters, making it a computationally efficient model. This section shows how we model the Transformer Gravity, combining the characteristics of the Gravity Model and the self-attention mechanism.
Problem Definition
Based on the pairs of source-destination ports obtained from the previous link prediction step, we encode the destination ports into 17 geographical regions according to the ISO-3166 standard [51], where ports within regions are expected to have a similar set of organisms and, therefore, share similar habitat. Over the encoded representation of regions, we now aim to estimate the sizes of shipping mobility flows between each source-destination pair $(P_{i},P_{j})$, where $P_{i}$ is the source port and $P_{j}$ is the destination port that pertains to a unique geographical region. A ship departing from a source port may have one or more destination ports in the same or different regions, and following the Deep Gravity [41] method, the goal is to estimate probabilities of the ships traveling to these geographical regions, becoming a multiclass classification task.
Predict target: We compile a set of 10 features from various sources for each pair $(P_{i},P_{j})$, such as shipping fluxes at ports, geodesic distances between source and destination, global economic trade volume, the graph metrics computed with the global shipping network, and others; detailed information about these extracted features is provided in Table 3 in the Methods. We represent the feature vector for each source-destination pair as $x_{ij}=\left<m_{1},m_{2},\ldots,m_{10}\right>$. Given the ships from each source port travel to multiple destination regions, these feature vectors are aggregated into a single data sample $X_{i}=\{x_{i1},x_{i2},\ldots,x_{iN}\}$, where $N$ is the number of destination regions in the data sample, and $1\leq N\leq 17$. Each destination region is represented as a class, so the prediction has $N$ classes for a sample. Since samples of varying lengths cannot be wrapped to a tensor for batch processing, we set the batch size to 1 for model input. Using $\hat{y}_{ij}$ as the estimated size of the mobility flow between $(P_{i},P_{j})$, which is the target of the prediction, we have:
where $O_{i}$ represents the total number of ships departing from source port $i$. $p_{ij}$ is the probability of ships traveling from source port $i$ to the destination region $j$, and $f(x_{ij})$ is the model output of feature vector $x_{ij}$.
Evaluation metric: The Common Part of Commuters (CPC) [52, 53, 54] is designed to measure the similarity between two sets of data, which could represent various aspects such as the volume of commuters, traffic, or trade between different locations. The metric calculates how much overlap there is between the predicted values $\hat{y}_{ij}$ and the actual values $y_{ij}$. In the context of Equation 4, $M$ represents the number of source ports and $N$ the number of destination regions. The values $\hat{y}_{ij}$ and $y_{ij}$ correspond to the flow of vessels from source port $i$ to destination region $j$, based on a model’s prediction and the actual observed values, respectively. A high value of the CPC, in this case, means that there is a large overlap between the predicted and actual datasets. Specifically, it would indicate that the predictions accurately capture the true data distribution patterns, with most predictive quantities closely matching the actual quantities. Contrarily, a low value of the CPC would suggest that there is little overlap between the predictions and the actual data, indicating that the model’s predictions diverge significantly from the observed data, which could be due to underprediction or overprediction in various parts or a general misalignment of the model with the reality. Accordingly, CPC considers the minimum common value between the predicted and actual data for each pair of source and destination ports, measuring the intersection over the values union:
For a better evaluation process, besides the CPC, we included the Normalized Root Mean Square Error ($NRMSE$) — lower is better — and Pearson Correlation Coefficients ($Corr.$) — higher is better — to measure the normalized errors and the correlation between the predictions and observations, defined as:
Model Framework
Our proposed Transformer Gravity model is composed of three main components: (1) the input embedding layer, which maps the input feature vectors to a higher-dimensional space that is compatible with the Transformer architecture; (2) the multilayer Transformer encoder, which involves the self-attention and feed-forward blocks that process the embeddings to capture complex relationships between input features; and, (3) the output linear layer, which maps the processed embeddings to the target flow predictions, computes loss and CPC and performs backpropagation based on the loss values. Figure 6 presents the model pipeline using two stacked Transformers modules and provides a glance at the layer’s relationships.
Figure 6: Framework of the Transformer Gravity model.
In our experiment, we define the number of heads as $h=2$, and the heads run operations in parallel.
Layer Normalization. After the multi-head attention layer, there is a dropout layer that randomly sets a certain percentage of elements to 0. The dropout ratio $p$ is set to $0.1$ as per Equation 9. Next, a skip connection is applied to add the input features $\textbf{z}_{0}$ to the output of the dropout layer $\textbf{z}_{dropout}$ before the self-attention block. This helps to retain the information from the input features and prevent vanishing gradients during backpropagation. The output of this connection, $\textbf{z}_{skip}$, is then normalized using Equation 9, where $\mu$ is the mean and $\sigma$ is the standard deviation of $\textbf{z}_{skip}$ with a small bias. The affine parameters $\alpha$ and $\beta$ are initialized as 1 and 0, respectively, and can be optimized during the training process.
Training and Optimization. Our Transformer Gravity model has three transformer encoder layers stacked together to capture complex input embedding dependencies, but the number of stacked layers can be changed to match different requirements and needs. The output value $\textbf{z}_{5}$ is obtained by passing the output of Equation 10 through a second and later third multi-head attention and feed-forward network block. The output value of the model is denoted as $f(x_{ij})$, and it produces a sequence of output values for a single data sample with a length of $N$. This sequence is then applied to a softmax function to produce probabilities $\{p_{i1},p_{i2},\ldots,p_{ij},\ldots,p_{iN}\}$ for $N$ classes. The predicted flow sizes $\{\hat{y}_{i1},\hat{y}_{i2},\ldots,\hat{y}_{ij},\ldots,\hat{y}_{iN}\}$ for each destination are obtained by multiplying these probabilities with the total outflows $O_{i}$ from the source port, as given in Equation 2. The loss for every sample is computed using Equation 3, and these losses are collected to derive the total loss. The model’s parameters are updated with each loss by processing a single sample. The summed CPC across all samples is calculated using Equation 4. After each training epoch, the summed CPC is divided by the number of samples $M$ (i.e., the number of source ports) to obtain the average CPC of that epoch. During training, we used the Adam optimizer with $L_{2}$ regularization on the weights and reduced the learning rate by a factor of 0.1 when there was no improvement with the validation CPC after 10 epochs. We applied early stopping when there was no improvement with the validation CPC after 20 epochs to prevent overfitting and improve training efficiency.
Experimental Results
During the experimentation, we used the proposed Transformer Gravity model with 1, 3, and 5 Transformer layers. This study compared the original 15-layer Deep Gravity model with a multi-layered variant of 3, 9, and 12 layers known as the Shallow Gravity models. Additionally, we experimented with machine learning regression models, such as linear-based, tree-based, and boosting-based models, as detailed in Table 2. The data from 2017 and 2018 formed the training data, while the 2019 data was used for testing.
@@
Transformer ${}^{\bigstar}$ 1 0.846 0.856 0.829 0.161 0.898 0.834 0.200 0.865 51,212
Transformer ${}^{\bigstar}$ 3 0.864 0.870 0.852 0.080 0.977 0.848 0.187 0.882 101,644
Transformer ${}^{\bigstar}$ 5 0.864 0.871 0.856 0.107 0.953 0.836 0.208 0.858 152,076
Table 2: Performance evaluation of the Transformer, Deep and Shallow Gravity models, and other baseline models. The cross-validation results present the mean ($CPC_{mean}$), maximum ($CPC_{max}$), minimum ($CPC_{min}$), mean Normalized Root Mean Square Error ($NRMSE$), and the mean Pearson Correlation Coefficients ($Corr.$) across five folds. Test results show the performance of the trained models on the entire test dataset. Details on the number of layers and parameters are provided for each deep learning model.
After analyzing $CPC$ values in Table 2, it can be observed that the Transformer Gravity models show a more stable performance than the competing models. The difference between the highest and lowest $CPC$ values for the Transformer Gravity models ranges from $0.015$ to $0.027$, which is lower than other gravity-based models ($0.019\sim 0.049$) and machine learning models ($0.164\sim 0.233$). This indicates that the variance in $CPC$ is lower when the flow prediction is done with the Transformer Gravity model. The variance can be attributed to the distribution of data samples across folds, as a random seed is set for the sample assignment. This way, the Transformer Gravity model’s performance is more consistent across different folds, meaning that given historical shipping data within the same temporal period, the Transformer Gravity model is more likely to predict shipping traffic flows with stable performance, regardless of the shipping traffic from which subsets of the data are sampled from.
@@
Table 2 shows that several baseline models have higher test $CPC$s than their best cross-validation $CPC$s. This difference in performance can be attributed to the distinct feature distributions in the test data compared to the cross-validation fold data. Additionally, the spatial and temporal dependencies intrinsic to the shipping data can also affect these evaluation results. Our training set includes shipping data from 2017 and 2018, while the test set comprises the entire year of 2019. Since both data sets represent complete years, the test data is expected to be more closely aligned with the training data in spatio-temporal distribution. This similarity can potentially explain the higher test performance than the validation results observed for these models. However, the evaluation of the Transformer Gravity model shows a better result in validation sets, which follow a more conventional pattern and have minimal variance over the cross-validation folds. This suggests that our model has learned more intrinsic connections from data features and is robust enough to overcome the impact potentially caused by the different distributions of datasets, demonstrating the model’s generalizability and reliability in predicting mobility flow patterns.
Table 2 also reveals that the performance of the Transformer Gravity model varies with the addition of layers. An increase in the number of the Transformer encoder layers initially enhances the model’s performance, as evidenced by the results of the 3-layer Transformer Gravity model compared with the single-layer model shown in the table. However, this performance increase is not unlimited. Comparing the 3-layer Transformer Gravity with the 5-layer one, as the model grows deeper with more integrated parameters, its performance tends to reach a plateau. Therefore, while stacking Transformers can be beneficial, it is important to optimize the number of layers to maximize the model’s performance.
@@
Risk Assessment of NIS through Ballast Water
As part of our study, we conducted a final experiment using the models we proposed on the global shipping network. Our goal was to assess the risk of NIS invasion associated with shipping flows using the BWRA decision tool, also employed by Transport Canada [11]. We aimed to demonstrate how our proposed model can improve real-world risk analysis, providing better information for regulating the oceans and making policies. The tool takes into consideration environmental conditions such as sea surface temperature and salinity at the locations where ballast water is taken and discharged. We gathered environmental variables, including minimum, maximum, and annual temperature, and annual salinity at these locations, and compiled them into a vector called $\nu_{i}=\langle t_{(i)_{min}},t_{(i)_{max}},t_{(i)},s_{(i)}\rangle$. Through the vector $\nu_{i}$, we proceed to calculate the environmental distance using the element-wise Euclidean distance calculation:
@@
Following the approach described above, we first integrated port data from World Port Index [46] with the environmental conditions in Global Port Environmental Data [55] to match the environmental variables with the ports. We then assigned these variables to each port involved in the ship movement of 2019, thus forming the environmental vector pairs $(\nu_{i},\nu_{j})$ for each pair of sources and destinations. Using Equation 11, we calculated the environmental distances for these pairs. In order to account for the differences in environmental distances across the shipping network, we included the frequency of each computed distance. This frequency is directly proportional to the predicted size of the shipping flow, which helps to adjust the environmental distance data to match the volume of shipping traffic. As illustrated in the pipeline diagram (Figure 5), we scaled the environmental distances using shipping volume from the Transformer Gravity model (denoted as $T(d)_{TG}$), the Deep Gravity model ($T(d)_{DG}$) and compared these with the scaled environmental distances based on actual shipping flows from 2019 (represented as $T(d)_{true}$). This approach was used to evaluate the dissimilarity of risk assessment results associated with the predicted and real shipping flows.
@@
Figure 7: Distribution of environmental distances for shipping flows in 2019. The true shipping flows are represented by the blue curve, while the dashed red and green curves depict the predictions from the Transformer Gravity (TG) and Deep Gravity (DG) models, respectively. The x-axis measures the environmental distance; smaller values indicate higher risk levels, and larger values indicate lower risks.
@@
Figure 7 shows the distributions of the three distance groups: $T(d)_{true}$, $T(d)_{TG}$, and $T(d)_{DG}$, from which we can find that $T(d)_{true}$ and $T(d)_{TG}$ are closely aligned in their trends, even in minor fluctuations. In contrast, $T(d)_{DG}$ is more differentiated from $T(d)_{true}$. To quantify the alignment between $T(d)_{TG}$ and $T(d)_{true}$, we calculated Pearson’s correlation coefficients for the two groups of environmental distances, reaching the value of $0.889$. This high coefficient indicates a strong linear correlation between the environmental distances associated with actual and predicted shipping flows. Figure 7 also reveals that the consistency between $T(d)_{TG}$ and $T(d)_{true}$ is more pronounced where the environmental distance is greater, implying a lower risk of invasion. Conversely, when smaller environmental distances indicate a higher invasion risk, the two groups show more discrepancies. Since the distance values in the two groups are calculated from the origin-destination pairs of the same year, these discrepancies can be attributed to the differences between the predicted and the real shipping flows. On the other hand, the Transformer Gravity model contains an important feature, i.e., exportation values, quantifying the annual trading volume between two countries in US dollars. When contributing to the model’s predictive performance, the feature value becomes zero and no longer positively associates trade with shipping flows when predicting inner-country shipping activities, and these are usually high-risk routes due to more environmental similarity between the source and destination locations. Further, there may also be critical factors in predicting high-risk shipping connections that remain unexplored, suggesting potential for further studies.
Although there are some discrepancies in the high-risk interval, the evaluation results show an overall low error of $0.208$ and a high similarity between the environmental distances scaled according to the predicted and actual shipping flows. This result suggests that the high performance of our Transformer Gravity model for ship traffic flow prediction contributes significantly to representing shipping intensity in BWRA. This has contributed to a more accurate risk assessment for the spread of NIS through ballast water and thus provides a valuable reference for global ballast water risk management in the future.

### v3

@@ Results
After analyzing the results from the model shown in Figure 4(a), we observed that the Random Forest model has a higher accuracy. However, the Logistic Regression model acts as a filter by removing some 1- and 2-trip trajectories before predicting ship traffic flow. These trajectories are considered unstable and unreliable for predicting future mobility flow and may be outliers. Therefore, the Logistic Regression model identifies and removes these trajectories, leaving behind only those predicted to have more stable traffic flow between origin-destination pairs. These trajectories are then utilized to collect features of the gravity-informed model for improving ship traffic flow prediction, such as seen in Figure 3. An illustrative data-flow representation of the Transformer Gravity model is depicted in Figure 5, while a comprehensive description and formalization of the model are thoroughly given within the Methods section.
Performance Evaluation of Ship Traffic Intensity
In this experiment, we aimed to evaluate and compare the performance of different models in predicting global ship traffic intensity. Our primary goal was to assess the effectiveness of our proposed Transformer Gravity model when compared with the original Deep Gravity model and various machine learning regression models. Through this comparative analysis, we aimed to identify the model(s) that best approximate real-world shipping mobility patterns and demonstrate superior predictive power. In deep learning models, the number of layers refers to the depth of the neural network. Each layer in the network performs operations on the input data, gradually transforming it to make predictions. Adding more layers allows the model to learn more complex patterns and relationships within the data. Therefore, the complexity and depth of the model architecture should be adjusted to see how it affects its performance in capturing and understanding the data. We used the proposed Transformer Gravity model with 1, 3, and 5 Transformer layers during the experimentation. This study compared the original 15-layer Deep Gravity model with a multi-layered variant of 3, 9, and 12 layers. Additionally, we experimented with machine learning regression models, such as linear-based, tree-based, and boosting-based models, as detailed in Table 1. The data from 2017 and 2018 formed the training data, while the 2019 data was used for testing.
@@
Transformer Gravity 1 0.846 0.856 0.829 0.161 0.898 0.834 0.200 0.865 51,212
Transformer Gravity 3 0.864 0.870 0.852 0.080 0.977 0.848 0.187 0.882 101,644
Transformer Gravity 5 0.864 0.871 0.856 0.107 0.953 0.836 0.208 0.858 152,076
Table 1: Performance evaluation of the Transformer, Deep Gravity models and their shallower-layered variants, and other baseline models. The cross-validation results present the mean ($CPC_{mean}$), maximum ($CPC_{max}$), minimum ($CPC_{min}$), mean Normalized Root Mean Square Error ($NRMSE$), and the mean Pearson Correlation Coefficients ($Corr.$) across five folds. Test results show the performance of the trained models on the entire test dataset. The number of layers and parameters are given for deep learning models.
Table 1 shows that the Transformer Gravity models show a more stable performance than the competing models. The difference between the highest and lowest $CPC$ values for the Transformer Gravity models ranges from $0.015$ to $0.027$, which is lower than other gravity-based models ($0.019\sim 0.049$) and machine learning models ($0.164\sim 0.233$). This indicates that the variance in $CPC$ is lower when the flow prediction is done with the Transformer Gravity model. The variance can be attributed to the distribution of data samples across folds, as a random seed is set for the sample assignment. This way, the Transformer Gravity model’s performance is more consistent across different folds, meaning that given historical shipping data within the same temporal period, the Transformer Gravity model is more likely to predict shipping traffic flows with stable performance, regardless of the shipping traffic from which data subsets are sampled.
@@
Table 1 shows that several baseline models have higher test $CPC$s than their best cross-validation $CPC$s. This difference in performance can be attributed to the distinct feature distributions in the test data compared to the cross-validation fold data. The spatial and temporal dependencies intrinsic to the shipping data can also affect these evaluation results. Our training set includes shipping data from 2017 and 2018, while the test set comprises 2019. Since both datasets represent complete years, the test data is expected to be more closely aligned with the spatial distribution of the training data. This may account for the discrepancy between the higher test performance and the observed validation results for these models. However, the evaluation of the Transformer Gravity model shows a better result in validation sets, which follow a more conventional pattern and have minimal variance over the cross-validation folds. This indicates that our model has gained a deeper understanding of the connections between data features and is robust enough to handle the impact of varying data distributions within the dataset. This shows the potential of the model to make accurate predictions on unseen datasets.
Table 1 also reveals that the performance of the Transformer Gravity model varies with the addition of layers. An increase in the number of the Transformer encoder layers initially enhances the model’s performance, as evidenced by the 3-layer Transformer Gravity model results compared with the single-layer model shown in the table. However, this performance increase is not unlimited. Comparing the 3-layer Transformer Gravity with the 5-layer one, its performance tends to plateau as the model grows deeper with better-optimized parameters. Therefore, while stacking Transformers can be beneficial, optimizing the number of layers is essential to maximize the model’s performance.
@@
Performance Evaluation over the Risk Assessment of Invasive Species
As part of our study, we conducted a final experiment using the models we proposed on the global shipping network. We assessed the risk of NIS introduction associated with shipping flows using the BWRA decision tool employed by Transport Canada [37, 38]. We aimed to demonstrate the performance of our proposed model over a real application scenario, whereby by having a more accurate overview of the shipping traffic between regions, we can improve real-world risk analysis. In this case, by forecasting shipping intensities on maritime pathways with varying levels of invasion threats, regulatory agencies can have access to an estimation that can aid data-driven ocean regulation and policy-making, especially by identifying high-risk pathways that need enhanced monitoring and regulations to support maritime biosecurity. The risk model we used produces an environmental distance value that compares the dissimilarities between a ship’s origin and destination. When a ship arrives at a destination and comes from an origin with similar environmental characteristics (i.e., higher) as the destination port, it is more likely to be inspected by the responsible authorities. The model considers environmental conditions such as sea surface temperature and salinity at the locations where ballast water is taken and discharged. We gathered environmental variables at these locations, including minimum, maximum, annual temperature, and yearly salinity. We compiled them into a vector called $\nu_{i}=\langle t_{(i)_{min}},t_{(i)_{max}},t_{(i)},s_{(i)}\rangle$. Through the vector $\nu_{i}$, we proceed to calculate the environmental distance using the element-wise Euclidean distance calculation:
@@
To evaluate the environmental distances between the shipping source and destination, we first integrated port data from World Port Index [36] with the environmental conditions in Global Port Environmental Data [39], which includes the annual average salinity, annual average, minimum and maximum temperature nearest to each port. For each port involved in the 2019 shipping data, we calculated the environmental distances based on the environmental conditions at both the source and destination ports using Equation 1. These environmental distance values inform the threat levels associated with various shipping pathways. Then, we integrated the environmental distance values with the shipping intensity by weighting environmental distance according to the number of trips on specific shipping routes. This combination of both threat levels and shipping intensity enables the mapping of the overall distribution of shipping activities across the global network at varying invasion risk levels. To assess the dissimilarity in risk assessment results between predicted and actual shipping flows, we examined the distributions of environmental distances using the estimated trips from the Transformer Gravity model (denoted as $T(d)_{TG}$), the Deep Gravity model ($T(d)_{DG}$), and the actual shipping flows from 2019 (represented as $T(d)_{true}$).
@@
Figure 6: Distribution of environmental distances for shipping flows in 2019. The blue curve represents the true shipping flows, while the dashed red and green curves depict the predictions from the Transformer Gravity (TG) and Deep Gravity (DG) models, respectively. The x-axis measures the environmental distance; smaller values indicate higher risk levels, and larger values indicate lower risks.
@@
Figure 6 shows the distributions of the three distance groups: $T(d)_{true}$, $T(d)_{TG}$, and $T(d)_{DG}$, from which we can find that $T(d)_{true}$ and $T(d)_{TG}$ are closely aligned in their trends, even in minor fluctuations. In contrast, $T(d)_{DG}$ is more differentiated from $T(d)_{true}$. To quantify the alignment between $T(d)_{TG}$ and $T(d)_{true}$, we calculated Pearson’s correlation coefficients for the two groups of environmental distances, reaching the value of $0.889$. This high coefficient indicates a strong linear correlation between the environmental distances associated with actual and predicted shipping flows. The figure also reveals that the consistency between $T(d)_{TG}$ and $T(d)_{true}$ is more pronounced where the environmental distance is greater, implying a lower risk of invasion. Conversely, the two groups show discrepancies in the higher invasion risk interval (i.e., smaller environmental distances), which can be attributed to the differences between the predicted and the actual shipping flows. These discrepancies can be influenced by an important feature, i.e., the bilateral trading volume, which quantifies the annual trading volume between two countries in US dollars. For inner-country shipping activities, the bilateral trading volume becomes zero and no longer positively associates trade with shipping flows. These inner-country routes are usually high-risk due to the environmental similarity between origins and destinations. Further, there may also be critical factors in predicting unexplored high-risk shipping connections, suggesting potential for further studies.